A total of 42 tons of two types of ore is to be added into a smelter. The first type contains 6% copper and the other contains 2.5% copper. Find the necessary amounts of each ore to produce 2 tons of copper.
I need two equations. Thank you so much

Answer by mananth(16946)

x 6% copper ore
y 2.5% copper ore
Total tons of ore
1.00 x + 1.00 y = 42.00 .............1
Quantity of copper
6.00 x + 2.50 y = 200.00 .............2
Eliminate y
multiply (1)by -2.50
Multiply (2) by 1.00
-2.50 x \ -2.50 y = -105.00
6.00 x + 2.50 y = 200.00
Add the two equations
3.50 x = 95.00
/ 3.50
x = 27.14
plug value of x in (1)
1.00 x + 1.00 y = 42.00
27.14 + y = 42.00
y = 42.00 -27
y = 14.86
y = 14.86
27.14 6% copper ore
14.86 2.5% copper ore
m.ananth@hotmail.ca